Circular No. 7867

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

XTE J1751–305 C. B. Markwardt, University of Maryland and Goddard Space Flight

Center (GSFC); and J. H. Swank, GSFC, report that RXTE PCA monitoring observations of the galactic-center region on Apr. 3.6 UT revealed a new millisecond pulsar, designated XTE J1751–305. The best fitting source position is α = 17^{h}51^{m}.5, δ = −30^{o}30′ (equinox 2000.0), with an estimated position uncertainty of ∼ 5′, due to nearby confusing sources (see item by in ’t Zand et al., below). In the first 200 s of a pointed RXTE PCA observation on Apr. 4.65, x-ray pulsations were detected (with 10σ significance) at a barycentered frequency of 435.313 Hz ± 0.005 Hz. The r.m.s. pulsed amplitude in the band 2–10 keV was 2.5 percent. The approximate 2–10-keV x-ray fluxes on Apr. 3.6 and 4.6 were 55 ± 1 and 60 ± 2 mCrab, respectively. More observations by RXTE are scheduled. This is only the second accretion-powered millisecond pulsar. Further observations at other wavelengths are encouraged.

J. J. M. in ’t Zand, Astronomical Institute, Utrecht University, and Space Research Organization Netherlands (SRON); P. Lowes, SRON and BeppoSAX Science Operation Center, Rome; J. Heise, SRON; and A. Bazzano and P. Ubertini, Istituto di Astrofisica Spaziale e Fisica Cosmica, Rome, report that XTE J1751–305 was detected in a BeppoSAX Wide Field Camera observation during Apr. 4.3–4.6 UT at a flux of 52 ± 3 mCrab (2–28 keV) and localized at α = 17^{h}51^{m}16^{s}, δ = −30^{o}37′.5 (equinox 2000.0; uncertainty 1′.2, 99-percent confidence). No systematic flux trends were observed at more than the 10-percent sensitivity, and no x-ray bursts were detected.

SUPERNOVA 2002br IN MCG +03-27-61 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2002br (cf. IAUC 7853), obtained by M. Calkins on Mar. 15.25 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a normal type-Ia supernova about a week after maximum. Adopting the NASA/IPAC Extragalactic Database recession velocity of 10 248 km/s for the host galaxy, the supernova expansion velocity is ∼ 10 200 km/s for Si II (rest 635.5 nm) and 13 200 km/s for Ca II (rest 395.1 nm).

2002 April 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green